OpenStudy (anonymous):

Suppose a moving object has 2000 J of kinetic energy. If its mass is 40 kg, what is its speed? a. 10 m/s b. 100 m/s c. 7.07 m/s d. 44.72 m/s

OpenStudy (jamesj):

What is the formula for kinetic energy?

OpenStudy (anonymous):

Ya. With the K.E formula, substitute the available values and find the non available value.

OpenStudy (anonymous):

use KE=1/2*m/v^2\[1/2 mv ^{2}\]

OpenStudy (anonymous):

ans wil be 10 m/s (a)

OpenStudy (anonymous):

The Formula for Kinetic Energy is 1/2 m v^2 So set 1/2 m v^2 = 2000 Joules Solve for v \[1/2mv^2=2000 J\]\[mv^2=4000 J\]\[v=_{4000/40}^{1/2}\] 10 m/s

OpenStudy (anonymous):

Ignore the v equation... I give up on the stupid thing for writing roots.